how many grams of neon will have the same numbere of atoms as 4gms of ca

how  many  grams  of neon  will have the same numbere of atoms as 4gms  of ca   

Grade:9

3 Answers

Sanju
6 years ago
Molar mass of Ne = 20g
Mass of 1 atom of Ne = Molar mass/Avagadro No = 20 / 6.022*10²³ g
Molar mass of Ca = 40 g
Number of particles in 4 g of Ca = 4/40*6.022*10²³ = 6.022*10²² atoms
Therefore, number of atoms of Ne same as 4 g of Ca = 20*6.022*10²²/6.022*10²³
=20/10
=2 g
Ans: 2g of Ne has the same no of atoms as 4g of Ca.
praveen kumar
6 years ago
no. of atoms is directly to the no. of moles
moles of Ca=4/40=1/10
moles of Ne=x/20=1/10
there fore x=2 g of the neon is necessary for the same no. of atoms in 4 g of Ca.
Shashank Gurunaga
6 years ago
atomic mass of Ne = 20 u or 20 g ….(Case 1)
moles = given Weight of substance/ Mass of the substance(molecular or atomic)
atomic mass of Ca = 40 u or 40g......(Case 2)
weight of each element given,
weight of Ne = x g
weight of Ca = 4g (from the question)
Using  the mole concept,
no of moles of Ne= no of moles of Ca
(using the mole formula)
x/20  =  4/40
from this x’s  values comes as 2g. Therefore weight of Ne = 2g.